OpenStudy (anonymous):

evaluate the integral integral of (ln(4x))^2dx

OpenStudy (anonymous):

im assuming this is integration by parts. but i am not sure what to make my u and what to make dv

OpenStudy (anonymous):

would ln(4x) be good for u? and i would use chain rule to find derivative?

OpenStudy (astrophysics):

\[\int\limits (\ln(4x))^2 dx\] is it like this?

OpenStudy (anonymous):

yes it is

OpenStudy (astrophysics):

You can just let u = (ln(4x))^2 and dv = 1

OpenStudy (anonymous):

hey astrophysics can you please help me I posted a question

OpenStudy (anonymous):

ok! where do you get the 1 from?

OpenStudy (anonymous):

pls help i need to get rest

OpenStudy (astrophysics):

We can think of there being a 1 when we have such functions dv=dx, it's like if you have to integrate tanx dx you would use by parts, lets u = tanx, and dv = dx

OpenStudy (astrophysics):

I guess it would've been better if I said dv = 1*dx

OpenStudy (anonymous):

ohh yes i see! thank you

OpenStudy (anonymous):

astro pls help

OpenStudy (astrophysics):

Not tanx I meant arctanx haha

OpenStudy (anonymous):

hmm im not getting the right answer.. is it correct to say that the integral of ln4 is = to xln4 ?

OpenStudy (anonymous):

\[(\ln(4x))^{2}x-\int\limits x \frac{ 2\ln4 }{ x } dx\]

OpenStudy (anonymous):

so then i cancel out the x on the outside and the x in the denominator..

OpenStudy (anonymous):

and i am left with the integral of 2ln4, so then i take out the 2 and put it behind the integral, and i am left with the integral of ln4

OpenStudy (anonymous):

from there i took the integral of ln4 which i believe is xln4 and i put both parts together but the answer is incorrect

OpenStudy (anonymous):

\[(\ln(4x))^2x-2xln4\] that is my answer..

OpenStudy (anonymous):

but the answer says incorrect..

imqwerty (imqwerty):

the answer is \[x(\ln (4x))^2-2xln (4x)+2x\] recheck once again :)

OpenStudy (jhannybean):

You're almost there.

OpenStudy (anonymous):

thank you.. i am not seeing where the 2x at the end comes from.. ?

OpenStudy (anonymous):

that is my work so far

imqwerty (imqwerty):

u=4x \[\int\limits\frac{ (\ln (u)) }{ 4}du \]after some vry bad calculations nd simplifications u get- \[\frac{ 1 }{ 4 }((u)(\ln (u))^2 -2\ln (u)+2(u))\]and then when u simplify nd put u=4x then u get that answer..

OpenStudy (anonymous):

ok i see, thank you!

imqwerty (imqwerty):

no prblm but that is a long method..